# WaveMamba: Spatial-Spectral Wavelet Mamba for Hyperspectral Image Classification

### v1

@@ III-B Experimental Results with CNN-based Models
Figure 2: The comparison above illustrates the accuracy and loss trends of CNN-based methods. The WaveFormer (represented by the red line) demonstrates quicker convergence than traditional CNN models.
@@
Figure 3: University of Houston: The proposed WaveMamba achieves OA=97.30% showing competitive performance. All these results are compiled using $10\times 10$ patch size with 25% training samples for all competing methods.
@@
Figure 4: Pavia University: The proposed WaveMamba achieves OA=showing competitive performance. All these results are compiled using $10\times 10$ patch size with 25% training samples for all competing methods.
TABLE II: Per Class classification accuracy (Overall (OA) and Average Accuracy (AA) %) and Kappa ($\kappa$) measure for University of Houston. (WM = WaveMamba).
Class [17] [18] [19] [23] [21] [22] [37] [24] [25] [27] [26] [20] [28] [29] [38] WM
1 93.31 97.26 93.39 97.24 83.00 88.13 81.23 80.51 79.47 93.44 93.48 95.36 92.67 90.90 92.88 97.92
2 97.81 97.29 99.54 97.69 83.74 90.04 87.59 67.45 79.08 95.37 95.10 98.58 97.57 95.11 95.99 98.72
4 100.0 95.78 98.09 97.85 92.57 89.49 77.05 97.39 98.51 97.69 97.13 97.82 96.58 92.27 99.17 98.39
5 98.16 98.41 97.70 99.91 99.81 98.20 90.09 96.64 100.0 98.61 97.66 99.99 99.78 98.58 98.29 99.67
6 100.00 91.36 100.0 92.46 100.0 99.30 93.94 84.66 87.53 95.08 97.37 98.42 97.47 96.42 96.39 93.82
7 87.83 94.60 90.96 90.70 87.59 98.97 87.53 78.78 77.00 91.46 91.93 84.32 86.43 89.18 91.60 87.85
8 85.91 91.82 89.18 96.78 73.22 98.01 93.02 45.49 49.77 78.02 94.88 80.80 83.25 83.90 83.33 97.42
9 75.33 92.39 90.62 94.85 81.21 95.94 86.12 76.05 72.19 92.26 88.57 78.82 82.61 86.58 92.05 91.69
10 82.52 90.61 93.22 99.27 68.15 98.65 78.68 36.93 94.51 97.61 89.76 95.10 96.50 99.09 98.05 98.20
11 79.19 89.09 87.91 96.13 89.85 87.38 81.65 64.05 85.59 93.82 95.44 95.82 93.46 92.34 92.48 96.27
12 72.76 94.18 83.15 99.18 89.15 94.33 89.71 70.01 59.63 92.45 93.15 89.65 89.55 91.27 91.24 98.70
13 79.49 82.51 84.11 65.16 92.28 98.95 85.49 98.81 80.06 95.86 82.75 97.75 92.98 91.22 92.87 82.90
OA 88.45 93.06 93.09 96.08 86.30 93.21 86.50 75.65 82.14 93.92 93.67 92.88 92.85 92.73 94.30 96.39
AA 87.81 86.61 92.06 94.93 88.68 93.22 85.25 79.34 84.19 94.65 94.03 94.10 93.79 93.51 94.96 96.07
$\kappa$ 87.50 92.50 92.53 95.76 85.14 92.63 85.85 73.73 80.70 93.43 93.16 92.30 92.27 92.14 93.84 96.10
TABLE III: Per Class classification accuracy (Overall (OA) and Average Accuracy (AA) %) and Kappa ($\kappa$) measure for Pavia University. (WM = WaveMamba).
Class [17] [18] [19] [23] [21] [22] [37] [24] [25] [27] [26] [20] [28] [29] [38] WM
1 92.67 95.21 95.25 96.34 90.44 94.75 98.76 78.09 76.94 97.06 98.74 86.16 87.16 90.59 95.70 96.92
2 92.79 92.54 96.64 99.51 96.63 95.54 95.46 84.67 93.70 92.29 99.51 9 4.50 95.05 94.62 94.05 99.58
3 90.60 91.18 81.74 82.74 93.50 99.83 87.75 53.47 81.10 99.97 90.81 93.94 93.95 94.39 99.61 87.70
4 98.15 97.21 95.50 96.33 90.94 93.99 91.37 75.09 78.42 97.12 92.74 88.97 92.76 84.61 98.92 95.88
5 98.28 100.0 99.59 100.00 98.71 95.82 92.13 99.83 100.0 100.0 99.53 98.95 98.85 99.18 100.0 99.40
6 93.29 99.93 93.43 92.31 98.82 97.32 97.10 52.40 64.24 99.40 90.96 96.07 99.19 99.53 99.19 98.40
7 83.01 95.75 89.24 92.06 98.13 96.58 96.47 78.74 96.27 100.0 93.80 99.58 99.08 99.34 99.93 97.89
8 84.50 97.59 88.81 90.67 93.95 95.37 86.25 84.35 58.58 98.99 89.83 86.63 91.74 97.10 98.50 88.86
9 99.77 99.47 99.53 93.30 95.24 92.18 89.27 99.60 98.16 99.47 96.82 95.54 95.49 93.95 99.96 95.56
OA 92.30 91.35 94.48 96.05 95.25 95.05 94.64 78.23 83.28 95.79 95.87 92.61 94.06 94.33 96.40 96.58
AA 88.86 85.07 93.15 93.71 95.15 93.01 92.78 78.47 83.06 98.25 94.75 93.37 94.81 94.81 98.43 97.11
$\kappa$ 89.66 88.94 92.67 94.74 93.55 94.70 93.54 71.55 77.74 94.54 94.58 90.29 92.22 92.60 95.31 96.17
III-C Comparative Results with Transformer and Mamba Networks
@@
To illustrate the effectiveness of WaveMamba, it was compared against a range of HSIC methods, including Spectralformer (SF) [17], HiT [18], CSiT [19], SSFTT [20], S3L [37], RIAN [21], CAT [22], SPRLT [24], CMT [25], and additional models such as DBDA [26], MSSG [27], LSFAT [28], CT-Mixer [29], SSFormer [23], and SS-Mamba [38]. This diverse set includes advanced CNN models incorporating dual-attention mechanisms, Transformer-based approaches, and hybrid models, thereby providing a comprehensive benchmark for assessing WaveMamba’s performance and innovation within HSIC techniques.
@@
The results, as presented in Tables III and II, demonstrate that WaveMamba surpasses the competing methods in terms of Overall Accuracy (OA), Average Accuracy (AA), and Kappa ($\kappa$) measure on both the Pavia University and University of Houston datasets. Specifically, WaveMamba achieves OA values of 96.58% and 96.39% for the respective datasets, outperforming the second-best method, SS-Mamba, by margins of 0.18% and 2.09%. Furthermore, WaveMamba exhibits superior performance in AA and $\kappa$ measure, underscoring its robustness and efficacy in HSIC tasks. The per-class classification accuracy results, detailed in Tables III and II, further emphasize WaveMamba’s strengths. Notable improvements are observed in classes 3, 6, 8, and 9 on the Pavia University dataset, and in classes 8, 10, 12, and 15 on the University of Houston dataset. These results highlight WaveMamba’s capability to accurately classify a variety of land cover types, affirming its potential as a promising approach for HSIC applications.
@@ IV Conclusions
This paper introduced ”WaveMamba,” which integrates wavelet transformation with the Spatial-Spectral Mamba architecture for HSIC. By leveraging wavelet-based extraction of multi-scale spatial-spectral features, WaveMamba captures both local texture patterns and global contextual relationships in an end-to-end trainable model. A key innovation is the incorporation of a State-Space Model, enhancing the modeling of temporal dependencies alongside spatial-spectral information. Extensive experiments show that WaveMamba significantly improves classification accuracy over well-known advanced DL methods, especially in datasets with limited training data, by effectively capturing nuanced spectral and structural details. The method not only achieves superior performance but also demonstrates robustness and generalizability, making it a promising solution for real-world HSIC challenges. Future work could focus on exploring self-supervised pre-training and further network optimizations to enhance WaveMamba’s performance in data-scarce environments.

### v2

@@ III-C Experimental Results with CNN-based Models
Figure 2: Accuracy and loss trends of CNN-based methods.
@@
Figure 3: The proposed WaveMamba achieves OA=97.30% showing competitive performance. All these results are compiled using $10\times 10$ patch size with 25% training samples for all competing methods.
@@
Figure 4: The proposed WaveMamba achieves OA=96.58% showing competitive performance. All these results are compiled using $10\times 10$ patch size with 25% training samples for all competing methods.
III-D Comparative Results with Transformer and Mamba
@@
To illustrate the effectiveness of WaveMamba, it was compared against a range of HSI classification methods, including Spectralformer (SF) [13], HiT [25], CSiT [26], SSFTT [33], S3L [28], RIAN [27], CAT [15], SPRLT [29], CMT [30], and additional models such as DBDA [32], MSSG [31], LSFAT [34], CT-Mixer [35], SSFormer [16], and SS-Mamba [36]. This diverse set includes advanced CNN models incorporating dual-attention mechanisms, Transformer-based approaches, and hybrid models, thereby providing a comprehensive benchmark for assessing WaveMamba’s performance.
@@
The results, as presented in Tables IV and III, demonstrate that WaveMamba surpasses the competing methods in terms of Overall Accuracy (OA), Average Accuracy (AA), and Kappa ($\kappa$) measure on both datasets. Specifically, WaveMamba achieves OA values of 97.80% and 98.63% for the respective datasets, outperforming the second-best method, SS-Mamba, by margins of 0.18% and 2.09%. Furthermore, WaveMamba exhibits superior performance in AA and $\kappa$ measure, underscoring its robustness and efficacy. Notable improvements are observed in classes 3, 6, 8, and 9 on the Pavia University dataset, and in classes 8, 10, 12, and 15 on the University of Houston dataset. These results highlight WaveMamba’s capability to accurately classify a variety of land cover types, affirming its potential as a promising approach.
@@ IV Conclusions
This paper introduced ”WaveMamba,” which integrates wavelet transformation with the Spatial-Spectral Mamba architecture for HSI classification. By leveraging wavelet-based extraction of multi-scale spatial-spectral features, WaveMamba captures both local texture patterns and global contextual relationships in an end-to-end trainable model. A key innovation is the incorporation of a State-Space Model, enhancing the modeling of temporal dependencies alongside spatial-spectral information. Extensive experiments show that WaveMamba significantly improves classification accuracy over well-known advanced DL methods, especially in datasets with limited training data, by effectively capturing nuanced spectral and structural details. The method not only achieves superior performance but also demonstrates robustness and generalizability, making it a promising solution for real-world HSI classification challenges. Future work could focus on exploring self-supervised pre-training and further network optimizations to enhance WaveMamba’s performance in data-scarce environments.